What is the molar concentration of As(III)

As(III) in a solution if 29.80 mL

mL of 0.100 M

M KBrO

3

KBrO3 is needed to titrate 49.15 mL

mL of the As(III)

As(III) solution? 


1
Expert's answer
2021-10-04T05:10:50-0400

"5Fe^2+MnO_4^-+8H^+=5Fe^3+Mn^2+4H_2O"

the reaction at titration

"\nu(MnO_4^-)=\nu(KMnO_4)=V(KMnO_4)*C(KMnO_4)"

"\nu(MnO_4^-)<<\nu(H^+)"

"\nu(MnO_4^-)=22,2 cm^3 * 0,0089 Moles/1000 cm^3 =0,00019758mol"

"5*\nu(MnO_4^-)=\nu(Fe^2+)=0,0009879moles"

"C(Fe^2+)=\nu(Fe^2+)/Voverall"

"Voverall = V(KMnO_4)+V(Fe^2\/\/Fe^3) + V(H_2SO_4)"

Voverall==22,2cm3+25cm3+20cm3=67,2 cm3

  1. C(Fe2+)=0,0147moles/dm3

1*. w(Fe2+)=C(Fe2+)*Molar/Dm3=0,8210g/dm3

w(Fe2+)/w(Fe3+)=0,821/(9,8-0,821)=0,09144


No specific salt anion is indicated (for example, it could be a complex like EDTA, which always binds to a metal ion in a 1 to 1 ratio).

Therefore, we will calculate based on the mass fractions calculated earlier.

(However, a calculation through equivalents would perhaps be more correct.)

(9,8-0,821)/9,8 * 100% = 91,62%